OpenStudy (anonymous):

Factors of x^4+x^2+1 are

OpenStudy (anonymous):

(x^2 + 1)^2

OpenStudy (anonymous):

But it's only x^2 not 2x^2

OpenStudy (anonymous):

(x^2 -x+1),(x^2 +x+1) are the factors

OpenStudy (anonymous):

Can you explain how?

OpenStudy (anonymous):

right! - sorry abdul i misread the question

OpenStudy (anonymous):

Ok, no problem @jimmyrep

OpenStudy (anonymous):

x^4 +x^2 +1 = x^4 +2x^2 +1 - x^2 = (x^2 +1)^2 - x^2 factoring it out would give answer

OpenStudy (anonymous):

Thanks ksram :D

OpenStudy (anonymous):

u r welcome :-)